### v4

# Excited $\Omega$ hyperon in charmful $\Omega_{b}$ weak decays

Abstract
We investigate the sextet $b$-baryon decay processes $\Omega_{b}\to J/\psi\Omega^{(*)}$, where $\Omega^{*}$ represents the $1P$-, $1D$- and $2S$-wave excited $\Omega$ hyperons in the spectroscopy. Using the constituent quark model, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega)=8.8\times 10^{-4}$, which agrees with the previous studies to the order of magnitude. By identifying $\Omega(2012)$ as $\Omega(1^{2}P_{3/2^{-}})$, ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$ can be similarly significant. Additionally, $\Omega(1^{2}D_{5/2^{+}})$ and $\Omega(1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}})$ states exhibit production rates of 0.5, and (0.6, 0.8), respectively, relative to their ground-state counterpart. Notably, our findings suggest that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))$ are as large as $(4.5,20)\times 10^{-4}$, making them accessible to experiments at LHCb.
I Introduction
In baryon spectroscopy, the formation of a colorless state from three quarks with three different colors is fundamentally linked to the principles of quantum chromodynamics (QCD) and hadron physics. While the combination of individual quark spins and orbital angular momenta is expected to produce a variety of light baryon states Klempt:2009pi ; Crede:2013kia ; Forkel:2008un , only a subset of these states has been observed to date pdg . The so-called “missing excited baryon” problem highlights an incomplete theoretical understanding of baryon spectroscopy Capstick:2000dk ; Klempt:2009pi .
The $\Omega$ hyperon spectroscopy has been relatively underexplored. Nonetheless, Belle initiated a new era of discoveries with the identification of an excited hyperon, $\Omega(2012)$ Belle:2018mqs , which was later reconfirmed through the process $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Additionally, a heavier excited hyperon $\Omega(2109)^{-}$, observed via $e^{+}e^{-}\to\Omega(2109)^{-}\bar{\Omega}^{+}+c.c.$, and reported by BESIII BESIII:2024eqk . To investigate the nature of these new excited hyperons, significant theoretical attention has been devoted Xiao:2018pwe ; Liu:2019wdr ; Wang:2018hmi ; Zeng:2020och ; Aliev:2018yjo ; Aliev:2018syi ; Arifi:2022ntc ; Polyakov:2018mow ; Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh . Notably, $\Omega(2012)$ is often preferred as an exotic molecule candidate Hu:2022pae ; Ikeno:2022jpe ; Lin:2019tex ; Lu:2020ste ; Ikeno:2020vqv ; Xie:2021dwe ; Valderrama:2018bmv ; Pavao:2018xub ; Huang:2018wth ; Gutsche:2019eoh rather than as a conventional $1P$-wave $\Omega(sss)$ state.
Clearly, the “missing excited baryon” problem persists in $\Omega$ hyperon spectroscopy, as all waves of $\Omega$ hyperons have been theoretically predicted Oh:2007cr ; Capstick:1986ter ; Faustov:2015eba ; Loring:2001ky ; Liu:2007yi ; Chao:1980em ; Chen:2009de ; An:2013zoa ; Kalman:1982ut ; Pervin:2007wa ; An:2014lga ; Engel:2013ig ; CLQCD:2015bgi ; Carlson:2000zr ; Goity:2003ab ; Schat:2001xr ; Matagne:2006zf ; Bijker:2000gq ; Aliev:2016jnp , but only a few have been experimentally observed. These include $\Omega\equiv\Omega(1672)$ Abrams:1964tu ; Barnes:1964pd , $\Omega(2012,2109)$ Belle:2018mqs ; Belle:2021gtf ; BESIII:2024eqk , and $\Omega(2250,2380,2470)$ Biagi:1985rn ; Aston:1987bb ; Aston:1988yn . Since further clarification and exploration are essential, we propose that recent measurements of the sextet $b$-baryon decay process $\Omega_{b}\to J/\psi\Omega$ D0:2008sbw ; CDF:2009sbo ; LHCb:2013wmn ; CDF:2014mon ; LHCb:2023qxn ; Nicolini:2023stq be extended to include $\Omega_{b}\to J/\psi\Omega^{*}$, where $\Omega^{*}$ denotes higher-wave excited $\Omega$ hyperons.
As LHCb continues to improve the statistics and precision LHCb:2023qxn ; Nicolini:2023stq , the absolute branching fractions of $\Omega_{b}^{-}$ decays may soon be measured with a similar accuracy to those of the anti-triplet $b$-baryon decays UA1:1991vse ; CDF:1992lrw ; CDF:2006eul ; CDF:1996rvy ; D0:2004quf ; D0:2007giz ; D0:2011pqa ; D0:2012hfl ; LHCb:2013hzx ; LHCb:2020iux ; LHCb:2019fim ; LHCb:2019aci ; ATLAS:2014swk ; ATLAS:2015hik ; CMS:2018wjk ; D0:2007gjs . Theoretical estimations are therefore needed, and a variety of theoretical tools are already available for such studies. These include the non-relativistic quark model Fayyazuddin:1998ap ; Mott:2011cx ; Cheng:1995fe ; Cheng:1996cs , the covariant confined quark model Gutsche:2018utw ; Gutsche:2017wag ; Gutsche:2013oea ; Gutsche:2015lea , the covariant oscillator quark model Mohanta:1998iu , the relativistic three-quark model Ivanov:1997hi ; Ivanov:1997ra , the light-front quark model Hsiao:2021mlp ; Wei:2009np ; Zhu:2018jet ; Wang:2024mjw , the perturbative QCD approach Rui:2023fiz ; Chou:2001bn , the generalized factorization approach Hsiao:2015cda ; Fayyazuddin:2017sxq ; Hsiao:2015txa , and SU(3) flavor analysis Dery:2020lbc .
In this work, we employ the constituent quark model Niu:2020gjw ; Niu:2025lgt ; Wang:2022zja ; Niu:2021qcc ; Niu:2020aoz ; Pervin:2006ie ; Pervin:2005ve to provide our estimations. As we will demonstrate, this model can be extended to study both the $\Omega_{b}\to J/\psi\Omega$ and $\Omega_{b}\to J/\psi\Omega^{*}$ decay processes, enabling a systematic analysis of the missing excited $\Omega$ states. This paper is organized as follows: In Sec. II, we apply the constituent quark model to two-body nonleptonic weak decays of $\Omega_{b}\to J/\psi\Omega^{(*)}$. In Sec. III, we present the numerical results. Finally, we provide our discussions and conclusion in Sec. IV.
II framework
As shown in Fig. 1, the recently observed sextet $b$-baryon decay, $\Omega_{b}\to J/\psi\Omega$, proceeds exclusively via the internal $W$-boson emission diagram Cheng:2021qpd , where $\Omega$ is strait-forwardly formed through the $\Omega_{b}\to\Omega$ transition. The single decay topology, extended to $\Omega_{b}\to J/\psi\Omega^{*}$, also facilitates the investigation of the $1P$-, $2S$- and $1D$-wave excited $\Omega^{*}$ states.
The quark-level effective Hamiltonian of $b\to c\bar{c}s$ weak transition induces the doubly charmful decay channels $\Omega_{b}\to J/\psi\Omega^{(*)}$, given by Buchalla:1995vs
$H_{W}$ $=$ $\frac{G_{F}}{\sqrt{2}}V_{bc}V_{cs}^{*}(C_{1}\mathcal{O}_{1}+C_{2}\mathcal{O}_{2})\,,$ (1)
where $G_{F}$ is the Fermi constant, $C_{1}$ and $C_{2}$ are the Wilson coefficients, and ($V_{bc}$, $V_{cs}$) are the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements. Additionally, ${\cal O}_{1,2}$ are the current-current operators, which are written as
$\mathcal{O}_{1}$ $=$ $\bar{\psi}_{\bar{s}_{\beta}}\gamma_{\mu}(1-\gamma_{5})\psi_{c_{\beta}}\bar{\psi}_{\bar{c}_{\alpha}}\gamma^{\mu}(1-\gamma_{5})\psi_{b_{\alpha}}\,,$
$\mathcal{O}_{2}$ $=$ $\bar{\psi}_{\bar{s}_{\alpha}}\gamma_{\mu}(1-\gamma_{5})\psi_{c_{\beta}}\bar{\psi}_{\bar{c}_{\beta}}\gamma^{\mu}(1-\gamma_{5})\psi_{b_{\alpha}}\,.$ (2)
Here, $\psi_{j_{\delta}}$ is the $j$th quark field, where $j$ can be $s$, $c$ or $b$, and $\delta=(\alpha,\beta)$ is the color index. We define the convention for the quark and antiquark fields
$\psi_{c}(x)=\int\frac{dp}{(2\pi)^{3/2}}\left(\frac{m}{p^{0}}\right)^{1/2}\sum_{s}\left[u_{s}(p)b_{s,c}(p)e^{-ip\cdot x}+v_{s}(p)d^{\dagger}_{s,c}(p)e^{ip\cdot x}\right],$ (3)
$\bar{\psi}_{c}(x)=\int\frac{dp}{(2\pi)^{3/2}}\left(\frac{m}{p^{0}}\right)^{1/2}\sum_{s}\left[\bar{u}_{s}(p)b^{\dagger}_{s,c}(p)e^{ip\cdot x}+\bar{v}_{s}(p)d_{s,c}(p)e^{-ip\cdot x}\right].$
where $c$ is color quantum number, and
$u_{s}(p)=\sqrt{\frac{E+m}{2m}}\left(\begin{array}[]{c}\varphi_{s}\cr\frac{\mathbf{\sigma}\cdot\emph{{p}}}{E+m}\varphi_{s}\end{array}\right),v_{s}(p)=\sqrt{\frac{E+m}{2m}}\left(\begin{array}[]{c}\frac{\mathbf{\sigma}\cdot\emph{{p}}}{E+m}\chi_{s}\cr\chi_{s}\end{array}\right).$ (8)
Here, $s$ represents spin labeling, so
$\sum_{s}\varphi_{s}\varphi_{s}^{{\dagger}}=\sum_{s}\chi_{s}\chi_{s}^{{\dagger}}=1.$ (9)
The anticommutation relations of the creation and annihilation operators are given by
$\{b_{s,c}(p),b_{s^{\prime},c^{\prime}}^{\dagger}(p^{\prime})\}=\{d_{s,c}(p),d_{s^{\prime},c^{\prime}}^{\dagger}(p^{\prime})\}=\delta_{ss^{\prime}}\delta_{cc^{\prime}}\delta^{3}(p-p^{\prime}).$ (10)
The normalization of spinor is
$\bar{u}_{s}(p)u_{s^{\prime}}(p)=-\bar{v}_{s}(p)v_{s^{\prime}}(p)=\delta_{ss^{\prime}}.$ (11)
Figure 1: Feynman diagram for the nonleptonic weak decay $\Omega_{b}\rightarrow J/\psi\Omega^{(*)}$.
According to Refs. Mannel:1992ti ; Buras:1985xv ; Blok:1992na , $C_{2}\langle J/\psi\Omega|{\cal O}_{2}|\Omega_{b}\rangle$ corresponds to the internal $W$-boson emission diagram. On the other hand, $C_{1}\langle J/\psi\Omega|{\cal O}_{1}|\Omega_{b}\rangle$ is replaced by $C_{1}/N_{c}^{eff}\langle J/\psi\Omega|{\cal O}_{2}|\Omega_{b}\rangle$ in the generalized factorization, where $N_{c}^{eff}$ is an effective color number that accounts for the non-factorizable QCD corrections. In the large $N_{c}^{eff}$ limit ($N_{c}^{eff}\to\infty$Mannel:1992ti ; Buras:1985xv ; Blok:1992na , the contribution from $C_{1}\mathcal{O}_{1}$ becomes negligible. Therefore, we only consider $C_{2}\mathcal{O}_{2}$ in this work.
The constituent quark model separates $H_{W}$ into the parity-conserving (PC) and parity-violating (PV) components Niu:2020gjw :
$H_{W}=H_{W}^{PC}+H_{W}^{PV}\,.$ (12)
In the non-relativistic approximation, the two components are given by Niu:2020gjw ; Niu:2025lgt ; Wang:2022zja
$H_{W}^{PC}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}C_{2}\frac{\hat{\phi}_{c}\hat{O}_{f}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})$
[…]
$-\left[\left(\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}+\frac{\textbf{p}_{3}}{2m_{3}}\right)-i\mbox{\boldmath$\sigma$\unboldmath}_{3}\times\left(\frac{\textbf{p}_{3}}{2m_{3}}-\frac{\textbf{p}_{3}^{\prime}}{2m_{3}^{\prime}}\right)\right]\cdot\mbox{\boldmath$\sigma$\unboldmath}_{4}$
[…]
$H_{W}^{PV}$ $\simeq$ $\frac{G_{F}}{\sqrt{2}}V_{cs}^{*}V_{cb}C_{2}\frac{\hat{\phi}_{c}\hat{O}^{f}}{(2\pi)^{3}}\delta^{3}(\textbf{p}_{3}-\textbf{p}_{3}^{\prime}-\textbf{p}_{4}-\textbf{p}_{5})$ (13)
[…]
Explicitly, we present
$\hat{O}_{f}$ $=$ $\hat{b}_{3^{\prime}}^{\dagger}(c)\hat{b}_{4}^{\dagger}(\bar{c})\hat{b}_{5}^{\dagger}(s)\hat{b}_{3}(b)\,,$
$\hat{\phi}_{c}$ $=$ $\delta_{c_{4}c_{3^{\prime}}}\delta_{c_{5}c_{3}}\,,$ (14)
where $\hat{b}^{({\dagger})}(q)$ annihilates (creates) the quark (q), and the $\delta_{c_{5}c_{3}}(\delta_{c_{4}c_{3^{\prime}}})$ represents that the colors of $c_{3}(c_{3^{\prime}})$ and $c_{5}(c_{4})$ should be conserved.
For the color operator to act on the color wave functions of $\Omega_{b}$, $\Omega^{(*)}$, and $J/\psi$, we define
$|\zeta_{\Omega_{b}}\rangle=\frac{1}{\sqrt{6}}(RGB-RBG+GBR-GRB+BRG-BGR)\,,$
$|\zeta_{\Omega^{(*)}}\rangle=|\zeta_{\Omega_{b}}\rangle\,,$
$|\zeta_{J/\psi}\rangle=\frac{1}{\sqrt{3}}(R\bar{R}+G\bar{G}+B\bar{B})\,.$ (15)
From Eqs. (II) and (II), we obtain
$\langle\zeta_{J/\psi}\zeta_{\Omega^{(*)}}|\hat{\phi}_{c}|\zeta_{\Omega_{b}}\rangle=\sqrt{3}\,.$ (16)
In the constituent quark model, the decay amplitude is represented as
$\mathcal{M}(\Omega_{b}\to J/\psi\Omega^{(*)})=\mathcal{M}^{PC}+\mathcal{M}^{PV}\,,$ (17)
where
$\mathcal{M}^{PC(PV)}$ $=$ $\frac{1}{\sqrt{3}}\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|$ (18)
$(1+\hat{P}_{25}+\hat{P}_{15})H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle\,.$
Here, the factor $\frac{1}{\sqrt{3}}$ is the normalization coefficient and $\hat{P}_{ij}$ represents the permutation operator of quarks $i$ and $j$. In the above equation, $J/\psi(\textbf{q};1,J_{J/\psi}^{z})$, $\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})$, and $\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})$ are wave functions, where ${\bf q}$ and ${\bf P}_{i(f)}$ are the total momentum, $J_{i(f)}$ are the total angular momentum, and $J^{z}_{i(f)}$ are the third component of the total angular momentum.
It is worth noting that the calculation of the amplitudes for $\Omega_{b}\to\Omega^{(*)}J/\psi$ decays should consider a question of symmetry. The flavor wave functions of $\Omega_{b}$ and $\Omega$ are $\varphi_{\Omega_{b}}=ssb$ and $\varphi_{\Omega}=sss$, respectively. In the $\Omega_{b}\rightarrow\Omega^{(*)}J/\psi$ nonleptonic decay process, the bottom quark of the $\Omega_{b}$ decays into an $s$ quark, which can be any of the three $s$ quarks in the $\Omega$. Since the wave functions of the final state $\Omega^{(*)}$ baryon are fully antisymmetric under the exchange of any two quarks, so we obtain
$\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle$
$=$ $\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|\hat{P}_{25}H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle$
$=$ $\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|\hat{P}_{15}H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle.$
Therefore, one can obtain
$\mathcal{M}^{PC(PV)}$ $=$ $\sqrt{3}\langle J/\psi(\textbf{q};1,J_{J/\psi}^{z})\Omega^{(*)}(\textbf{P}_{f};J_{f},J_{f}^{z})|$ (20)
$H_{W}^{PC(PV)}|\Omega_{b}(\textbf{P}_{i};J_{i},J_{i}^{z})\rangle.$
Similar calculations have been applied in Ref. Pervin:2006ie .
Table 1: The spin-flavor-space wave-functions of $\Omega$ baryons under $SU(6)$ quark model classification are listed below. We denote the baryon states as $|N_{6},^{2S+1}N_{3},N,L,J^{P}\rangle$ where $N_{6}$ stands for the irreducible representation of spin-flavor $SU(6)$ group, $N_{3}$ stands for the irreducible representation of flavor $SU(3)$ group and $N$, $L$, $J^{P}$ as principal quantum number, total orbital angular momentum and spin-parity, respectivelyXiao:2018pwe ; Liu:2019wdr ; Wang:2018hmi . The $\phi$, $\chi$, $\psi$ denote flavor, spin and spatial wave function, respectively. The Clebsch-Gorden coefficients of spin-orbital coupling and color wave function $\zeta_{\Omega}$ have been omitted.
$n^{2S+1}L_{J^{p}}$ $|N_{6},^{2S+1}N_{3},N,L,J^{P}\rangle$ Wave function
$1^{4}S_{\frac{3}{2}^{+}}$ $|56,^{4}10,0,0,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{000}^{s}$
$1^{2}P_{\frac{1}{2}^{-}}$ $|70,^{2}10,1,1,\frac{1}{2}^{-}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{11L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{11L_{z}}^{\lambda}\right)$
$1^{2}P_{\frac{3}{2}^{-}}$ $|70,^{2}10,1,1,\frac{3}{2}^{-}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{11L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{11L_{z}}^{\lambda}\right)$
$2^{2}S_{\frac{1}{2}^{+}}$ $|70,^{2}10,2,0,\frac{1}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{200}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{200}^{\lambda}\right)$
$2^{4}S_{\frac{3}{2}^{+}}$ $|56,^{4}10,2,0,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{200}^{s}$
$1^{2}D_{\frac{3}{2}^{+}}$ $|70,^{2}10,2,2,\frac{3}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{22L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{22L_{z}}^{\lambda}\right)$
$1^{2}D_{\frac{5}{2}^{+}}$ $|70,^{2}10,2,2,\frac{5}{2}^{+}\rangle$ $\frac{1}{\sqrt{2}}\left(\phi^{s}\chi^{\rho}\psi_{22L_{z}}^{\rho}+\phi^{s}\chi^{\lambda}\psi_{22L_{z}}^{\lambda}\right)$
$1^{4}D_{\frac{1}{2}^{+}}$ $|56,^{4}10,2,2,\frac{1}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{3}{2}^{+}}$ $|56,^{4}10,2,2,\frac{3}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{5}{2}^{+}}$ $|56,^{4}10,2,2,\frac{5}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
$1^{4}D_{\frac{7}{2}^{+}}$ $|56,^{4}10,2,2,\frac{7}{2}^{+}\rangle$ $\phi^{s}\chi^{s}\psi_{22L_{z}}^{s}$
Table 2: The spatial functions $\psi^{\sigma}_{NLM_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$ as the linear combination of $\psi_{n_{\rho}l_{\rho}m_{\rho}}(\mathbf{p}_{\rho})\psi_{n_{\lambda}l_{\lambda}m_{\lambda}}(\mathbf{p}_{\lambda})$.
$\psi^{S}_{000}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$
$\psi^{\rho}_{11M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{01M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$
$\psi^{\lambda}_{11M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{01M_{L}}(\mathbf{p}_{\lambda})$
$\psi^{S}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{100}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{100}(\mathbf{p}_{\lambda})]$
$\psi^{\lambda}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[-\psi_{100}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{100}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{200}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$
$=\frac{1}{\sqrt{3}}[\psi_{011}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})-\psi_{010}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{01-1}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})]$
$\psi^{S}_{22M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{02M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})+\psi_{000}(\mathbf{p}_{\rho})\psi_{02M_{L}}(\mathbf{p}_{\lambda})]$
$\psi^{\lambda}_{22M_{L}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{02M_{L}}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})-\psi_{000}(\mathbf{p}_{\rho})\psi_{02M_{L}}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{222}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{011}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})$
$\psi^{\rho}_{221}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{010}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})+\psi_{011}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{220}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})$
$=\frac{1}{\sqrt{6}}[\psi_{01-1}(\mathbf{p}_{\rho})\psi_{011}(\mathbf{p}_{\lambda})+2\psi_{010}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{011}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{22-1}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\frac{1}{\sqrt{2}}[\psi_{01-1}(\mathbf{p}_{\rho})\psi_{010}(\mathbf{p}_{\lambda})+\psi_{010}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})]$
$\psi^{\rho}_{22-2}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{01-1}(\mathbf{p}_{\rho})\psi_{01-1}(\mathbf{p}_{\lambda})$
To calculate $\mathcal{M}^{PC(PV)}$ in Eq. (20), the details of the wave functions are required. The total wave function of a baryon system should include four parts: a color wave function $\zeta$, a flavor wave function $\phi$, a spin wave function $\chi$, and a spatial wave function $\psi$. The spin wave function of a baryon satisfies SU(2) symmetry, which can be expressed as symmetric ($\chi^{S}_{S_{z}}$), mixed antisymmetric($\chi^{\rho}_{S_{z}}$), mixed symmetric($\chi^{\lambda}_{S_{z}}$) spin wave functions. The detailed representations of the spin wave function is shown in Refs. Niu:2020gjw ; Wang:2017kfr ; Xiao:2013xi .
For the $\Omega$ spectrum, the flavor-spin wave functions are representations of SU(6), which are denoted by $|N_{6},^{2S+1}N_{3}\rangle$, where $N_{6}$ ($N_{3}$) represents the SU(6) (SU(3)) representation and $S$ stands for the total spin of the wave function, and the detailed wave function of $\Omega^{(*)}$ excited states are shown in Table 1. In momentum space, we present the simple harmonic oscillator wave functions to describe the $\Omega^{(*)}$ baryon Wang:2018hmi . The explicit forms of the spatial wave function $\Psi_{NLM_{L}}^{\sigma}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$ in the momentum space, up to the $N$ = 2 shell, can be found in Table 2. The superscript $\sigma$ characterizes the same set of quantum numbers $(N,L,M_{L})$ arising from different combinations of the $\rho$ and $\lambda$ oscillation systems Karl:1969iz ; Zhenping:1991 ; Xiao:2013xi . Specifically, the combinations are defined by $N=2(n_{\rho}+n_{\lambda})+l_{\rho}+l_{\lambda}$, $L=l_{\rho}+l_{\lambda},...,|l_{\rho}-l_{\lambda}|$, and $M_{L}=m_{\rho}+m_{\lambda}$, where $(n_{i},l_{i},m_{i}$) with $i=\rho$ or $\lambda$ are the principal, orbital, and magnetic quantum numbers, respectively. The internal momenta of the $\rho$- and $\lambda$-oscillators, $\textbf{p}_{\rho}$ and $\textbf{p}_{\lambda}$, are expressed as
$\mathbf{p}_{\rho}$ $=$ $\frac{\sqrt{2}}{2}(\textbf{p}_{1}-\textbf{p}_{2})\,,$
$\mathbf{p}_{\lambda}$ $=$ $\frac{\sqrt{6}}{2}\frac{m_{5}(\textbf{p}_{1}+\textbf{p}_{2})-(m_{1}+m_{2})\textbf{p}_{5}}{m_{1}+m_{2}+m_{5}}\,,$ (21)
respectively.
In momentum space, $\psi_{nlm}(\textbf{p})$ is expressed by
$\psi_{nlm}(\textbf{p})=(i)^{l}(-1)^{n}\left[\frac{2n!}{(n+l+1/2)!}\right]^{1/2}\frac{1}{\alpha^{l+3/2}}$
$\mathrm{exp}\left(-\frac{\textbf{p}^{2}}{2\alpha^{2}}\right)L_{n}^{l+1/2}(\textbf{p}^{2}/\alpha^{2})\mathcal{Y}_{lm}(\textbf{p})\,,$ (22)
where the $l$th solid harmonic polynomial is defined as $\mathcal{Y}_{lm}(\textbf{p})=|\textbf{p}|^{l}Y_{lm}(\mathbf{\hat{p}})$, and the oscillator parameter $\alpha$ can be either $\alpha_{\rho}$ or $\alpha_{\lambda}$. For the $\Omega^{(*)}$ wave function, we define $\alpha\equiv\alpha_{\rho}=\alpha_{\lambda}$. In this work, we use the $\alpha$ value obtained from single Gaussian fitting in Ref. Liu:2019wdr .
For $\Omega_{b}$ baryon, the total wave function can be represented as
$\Psi_{\Omega_{b}}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\zeta_{\Omega_{b}}\varphi_{\Omega_{b}}\chi^{\lambda}_{s_{z}}\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})\,,$ (23)
where $\zeta_{\Omega_{b}}$, $\chi^{\lambda}_{S_{z}}$, and $\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$ represent the color, spin, and spatial wave functions, respectively. It is worth noting that in momentum space, we employ the simple harmonic oscillator wave functions to describe $\psi_{000}^{S}(\textbf{p}_{\rho},\textbf{p}_{\lambda})$, which is represented as
$\psi^{S}_{000}(\mathbf{p}_{\rho},\mathbf{p}_{\lambda})=\psi_{000}(\mathbf{p}_{\rho})\psi_{000}(\mathbf{p}_{\lambda})$ (24)
where $\textbf{p}_{\rho}$ and $\textbf{p}_{\lambda}$ are the internal momenta of the $\rho$- and $\lambda$-oscillators, respectively. It is worth mentioning that the parameter $\alpha_{\lambda}$ of $\Omega_{b}$ wave function is given by $\alpha_{\lambda}=[(3m_{b})/(2m_{s}+m_{b})]^{1/4}\alpha_{\rho}$ Wang:2017kfr ; Yao:2018jmc .
For the wave function of $J/\psi$ meson, which can be represented as
$\Psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\zeta_{J/\psi}\varphi_{J/\psi}\chi^{1}_{s_{z}}\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})\,,$ (25)
where $\zeta_{J/\psi}$ is the color wave function. $\varphi_{J/\psi}$, $\chi^{1}_{s_{z}}$, and $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ represent the flavor, spin, and spatial wave functions, respectively, which are given as:
$\chi^{1}_{1,0,-1}=(\uparrow\uparrow,(\uparrow\downarrow+\downarrow\uparrow)/\sqrt{2},\downarrow\downarrow)\,,$ (26)
$\varphi_{J/\psi}=c\bar{c}\,,$
$\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})=\frac{1}{\pi^{3/4}\beta^{3/2}}\mathrm{exp}\left[-\frac{(\textbf{p}^{\prime}_{3}-\textbf{p}_{4})^{2}}{8\beta^{2}}\right]\,.$
In Eq. (26), $s_{z}$ denotes the $z$ component of the $J/\psi$ spin and $\beta$ in $\psi(\textbf{p}^{\prime}_{3},\textbf{p}_{4})$ controls the width of Gaussian distribution in momentum space.
To convert ${\cal M}(\Omega_{b}\to J/\psi\Omega^{(*)})$ from Eqs. (17) and (18) into the decay width, we use the following equation:
$\Gamma=8\pi^{2}\frac{|\textbf{q}|E_{J/\psi}E_{\Omega^{(*)}}}{M_{\Omega_{b}}}\frac{1}{2J_{\Omega_{b}}+1}\sum_{J_{i}^{z},J_{f}^{z}}|\mathcal{M}_{J_{i}^{z},J_{f}^{z}}|^{2}\,,$ (27)
which is applied in our numerical analysis. In Eq. (27), $\mathcal{M}_{J_{i}^{z},J_{f}^{z}}$ is the transition amplitude.
III Numerical results
For our numerical analysis, we adopt the CKM matrix elements and the masses of $\Omega_{b}$ and $J/\psi$ from PDG pdg , while the quark masses are taken from Wang:2017kfr , as follows:
$(V_{cb},V_{cs})=(0.042,0.987)\,,$
$(m_{\Omega_{b}},m_{\Omega},m_{J/\psi})=(6.046,1.672,3.0969)~{}{\rm GeV}\,,$
$(m_{s},m_{c},m_{b})=(0.45,1.48,5.0)~{}{\rm GeV}\,.$ (28)
The lifetime of the $\Omega_{b}$ state is also taken from PDG pdg : $\tau_{\Omega_{b}}=1.64\times 10^{-12}s$. The Wilson coefficient $C_{2}=-0.365$ is from Hsiao:2021mlp ; Ali:1998eb ; Hsiao:2014mua . For the harmonic oscillator parameters, we use $(\alpha_{\lambda},\alpha_{\rho})=(0.56,0.44)$ GeV for $\Omega_{b}$ Zhong:2007gp , and $\beta=0.50$ GeV for $J/\psi$, as adopted in Barnes:2005pb ; Xiao:2018iez . For the ground-state and excited $\Omega$ hyperon states, the masses and the oscillator parameters $\alpha$, calculated in Liu:2019wdr , are summarized in Table 3.
Table 3: Our results for $\Omega_{b}\to J/\psi\Omega^{(*)}$ using the constituent model, with the quantum numbers $n^{2S+1}J^{P}$ assigned to $\Omega^{(*)}$. The parameter $\alpha$ and the mass of $\Omega^{(*)}$ $M_{f}$ are both given in units of MeV; $\Gamma$ and ${\cal B}$ are in units of $10^{-17}$ GeV and $10^{-4}$, respectively.
[…]
$\Omega(1^{4}S_{\frac{3}{2}^{+}})$ 440 1672 36 8.8 1.00
$\Omega(1^{2}P_{\frac{1}{2}^{-}})$ 428 1957 17 4.2 0.48
$\Omega(1^{2}P_{\frac{3}{2}^{-}})$ 411 2012 42 11 1.25
$\Omega(2^{2}S_{\frac{1}{2}^{+}})$ 387 2232 18 4.5 0.51
$\Omega(2^{4}S_{\frac{3}{2}^{+}})$ 381 2159 79 20 2.3
$\Omega(1^{2}D_{\frac{3}{2}^{+}})$ 394 2245 7.5 1.9 0.22
$\Omega(1^{2}D_{\frac{5}{2}^{+}})$ 380 2303 17 4.2 0.48
$\Omega(1^{4}D_{\frac{1}{2}^{+}})$ 413 2141 9.3 2.3 0.26
$\Omega(1^{4}D_{\frac{3}{2}^{+}})$ 399 2188 21 5.4 0.61
$\Omega(1^{4}D_{\frac{5}{2}^{+}})$ 383 2252 29 7.2 0.81
$\Omega(1^{4}D_{\frac{7}{2}^{+}})$ 367 2321 14 3.3 0.38
Utilizing the inputs above, we calculate ${\cal B}(\Omega_{b}\to J/\psi\Omega)$ to compare it with the previous studies. Specifically, the decay widths and branching fractions of possible $\Omega_{b}\to J/\psi\Omega^{(*)}$ are presented in Table 3. Additionally, we plot $\Gamma(\Omega_{b}\to J/\psi\Omega)$ as functions of the parameters $\alpha$ and $\beta$ in Fig. 2 and $\Gamma(\Omega_{b}\to J/\psi\Omega^{*})$ as a function of $m_{\Omega^{*}}$ in Fig. 3.
Table 4: The branching fraction of $\Omega_{b}\rightarrow J/\psi\Omega$ is expressed in units of $10^{-4}$, compared with results from different models.
Our work   Gutsche:2018utw   Fayyazuddin:2017sxq   Hsiao:2021mlp   Cheng:1996cs   Rui:2023fiz   Wang:2024mjw
[…]
Figure 2: The decay width of $\Omega_{b}\to J/\psi\Omega$ varies with the harmonic parameters $\alpha$ of $\Omega$ baryon and $\beta$ of $J/\psi$ meson.
IV Discussions and conclusion
The branching fraction of $\Omega_{b}\to J/\psi\Omega$, on the order of $10^{-5}-10^{-3}$, indicates inconclusive calculations from Gutsche:2018utw ; Fayyazuddin:2017sxq ; Hsiao:2021mlp ; Cheng:1996cs ; Rui:2023fiz . On the other hand, the constituent quark model has been applied to $\Omega_{c}^{0}$ decays Wang:2022zja , where ${\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-})/{\cal B}(\Omega_{c}^{0}\to\pi^{+}\Omega^{-})=0.21$ agrees with the experimental value of $0.220\pm 0.059\pm 0.035$ Belle:2021gtf , suggesting its applicability to $\Omega_{b}$ decays.
[…]
It should be pointed out that our calculation of $\Omega_{b}\rightarrow\Omega J/\psi$ depends on $\alpha$ and $\beta$. To test the sensitivity, we depict $\Gamma(\Omega_{b}\rightarrow J/\psi\Omega)$ as a function of $\alpha$ and $\beta$, within the parameter space: $0.4~{}{\rm GeV}<\alpha<0.5~{}{\rm GeV}$ and $0.48~{}{\rm GeV}<\beta<0.55~{}{\rm GeV}$. In Fig. 2, it can be seen that $\Gamma(\Omega_{b}\rightarrow J/\psi\Omega)$ is more sensitive to $\beta$, varying from 3.0 to 4.2 times $10^{-16}$ GeV. However, it remains close to the central value: $3.6\times 10^{-16}$ GeV (see Table 3).
There are two excited $1P$-wave $\Omega$ states in the $\Omega$ hyperon spectroscopy: $\Omega(1^{2}P_{1/2^{-}})$ and $\Omega(1^{2}P_{3/2^{-}})$. It is interesting to note that the newly observed $\Omega(2012)$ state is more likely to be assigned to the latter: $\Omega(1^{2}P_{3/2^{-}})$, based on the fact that the measured mass and decay width are consistent with the quark model predictions Xiao:2018pwe ; Liu:2019wdr ; Aliev:2018yjo ; Aliev:2018syi ; Polyakov:2018mow .
Based on this assignment, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$. Specifically, ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))/{\cal B}(\Omega_{b}\to J/\psi\Omega)\simeq 1.3$, which clearly avoids the uncertain fragmentation fraction $f_{\Omega_{b}}$, making it beneficial for experimental examination. In fact, $\Omega(2012)$ has been observed Belle:2018mqs and reconfirmed in $\Omega_{c}^{0}\to\pi^{+}\Omega(2012)^{-},\Omega(2012)^{-}\to\Xi^{0}K^{-}$ Belle:2021gtf . Likewise, this can be reconfirmed from the resonant $\Omega_{b}$ decays, such as $\Omega_{b}\to J/\psi\Omega(2012)$, followed by the subsequent decay $\Omega(2012)\rightarrow\Xi^{0}K^{-}$. In particular, $\Omega(2012)\to\Xi K$ has a branching fraction of around $90\%$ Zhong:2022cjx . The $\Omega(1^{2}P_{1/2^{-}})$ is a typically missing $\Omega$ hyperon, with a predicted mass around 1950 MeV Engel:2013ig ; Capstick:1986ter ; Faustov:2015eba ; Liu:2019wdr . In our evaluation, ${\cal B}(\Omega_{b}\to J/\psi\Omega(1^{2}P_{1/2^{-}}))=4.2\times 10^{-4}$ indicates that $\Omega(1^{2}P_{1/2^{-}})$ has a smaller but compatible production rate to its $1P$-wave cousin state.
Figure 3: The decay widths of $\Omega_{b}\rightarrow J/\psi\Omega^{*}$ as a function of $M_{\Omega^{*}}$, with the range from the theoretical estimations.
In the quark model, the $1D$-wave $\Omega^{*}$ hyperons are classified into the spin-doublet $\Omega(1^{2}D_{3/2^{+}},1^{2}D_{5/2^{+}})$ and the spin-quartet $\Omega(1^{4}D_{1/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$, with masses evaluated to be around 2150 to 2250 MeV Liu:2019wdr ; Xiao:2018pwe ; Wang:2018hmi ; Wang:2022zja in the mass spectrum. However, these states have yet to be firmly established experimentally.
In accordance with the masses from Ref. Liu:2019wdr , we calculate $\Omega_{b}\to J/\psi\Omega(1D)$. Our the results are listed in Table 3, where $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have production rates to be 0.5, 0.6, 0.8, 0.4, respectively, relative to their ground-state counterpart. We highlight that ${\cal B}(\Omega_{b}\to J/\psi\Omega(1^{4}D_{5/2^{+}}))=7.2\times 10^{-4}$ is sufficiently large for a promising establishment of $\Omega(1^{4}D_{5/2^{+}})$ through $\Omega_{b}\to\Omega(1^{4}D_{5/2^{+}})J/\psi$, followed by $\Omega(1^{4}D_{5/2^{+}})\to\Xi(1530)^{0}K^{-}$, where ${\cal B}(\Omega(1^{4}D_{5/2^{+}})\to\Xi(1530)K)$ is as large as $80\%$ Liu:2019wdr ; Xiao:2018pwe .
As for the existing $2S$-wave $\Omega^{*}$ hyperon states, $\Omega(2^{2}S_{1/2^{+}})$ and $\Omega(2^{4}S_{3/2^{+}})$, we obtain ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))=(4.5,20)\times 10^{-4}$. Remarkably, in the current $\Omega$ spectroscopy studies, $\Omega(2^{4}S_{3/2^{+}})$ in $\Omega_{b}\to J/\psi\Omega^{*}$ is the only one that can have a production rate larger than that of its ground state counterpart.
The constituent quark model relies on theoretical inputs from the $\Omega^{*}$ masses. Due to the nature of these theoretical estimations, the masses carry some uncertainties. Therefore, in Figs. 3, we depict the decay widths involving various excited $\Omega$ hyperons as functions of $M_{\Omega^{*}}$, with the possible ranges determined by the model calculations. It is found that the predicted decay widths for both $\Omega^{*}(1P,1D)$ and $\Omega^{*}(2S)$ are not sensitive to the uncertainties arising from the mass estimations.
In conclusion, we have calculated the sextet $b$-baryon decays $\Omega_{b}\to J/\psi\Omega^{(*)}$ using the constituent quark model. We found that ${\cal B}(\Omega_{b}\to J/\psi\Omega)=8.8\times 10^{-4}$, which is consistent with the previous studies. With $\Omega(2012)$ identified as $\Omega(1^{2}P_{3/2^{-}})$, we found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2012))=1.1\times 10^{-3}$, which is compatible with ${\cal B}(\Omega_{b}\to J/\psi\Omega)$. Additionally, the production rates of the $\Omega(1^{2}D_{5/2^{+}},1^{4}D_{3/2^{+}},1^{4}D_{5/2^{+}},1^{4}D_{7/2^{+}})$ states have been calculated to be 0.5, 0.6, 0.8, 0.4, respectively, relative to their ground-state counterpart. We also found that ${\cal B}(\Omega_{b}\to J/\psi\Omega(2^{2}S_{1/2^{+}},2^{4}S_{3/2^{+}}))=(4.5,20)\times 10^{-4}$, which is promising for measurement by LHCb. Since we have demonstrated that our calculations are insensitive to the parameter inputs and uncertainties arising from the $\Omega^{*}$ masses, this provides a suitable test-bed to investigate the $\Omega$ hyperon spectroscopy.
Acknowledgements
We are very grateful to the referee for pointing out the incorrect calculation of the color factor, and to Qiang Zhao and Di Wang for their very helpful discussions regarding the color factor. This work was supported by the National Natural Science Foundation of China (Grants No.12205026, No.12175065 and No.12235018), and Applied Basic Research Program of Shanxi Province, China under Grant No. 202103021223376. YKH was supported by NSFC (Grants No. 12175128 and No. 11675030). JW was supported by General Programs of Changzhi College.